OpenStudy (anonymous):

Determine whether or not the triangle with vertices a(1,2) , b(5,5) and c(-1,13) . Use graphical evidence to determine whether or not the triangle is a right triangle

OpenStudy (tkhunny):

1) Are they collinear? If not, they form a unique triangle.

OpenStudy (wolf1728):

Anyway I made a graphic It appears it is a right triangle.

OpenStudy (tkhunny):

"Appears"? It is or it isn't. Prove it!!

OpenStudy (wolf1728):

That's what I was going to work on next. Is there a quick test to determine a right angle OR do we have to determine lengths of each side?

OpenStudy (tkhunny):

Slopes of the intersecting lines?

OpenStudy (wolf1728):

slope (5,5) (1,2) = (5 -1)/(5 - 2) = 4/3 slope (-1, 13) (5,5) = (-1 -5) / (13 -5) = -6 / 8 = -3/4

OpenStudy (wolf1728):

slopes are 90° different therefore it is a right triangle.

OpenStudy (tkhunny):

Very nice.

OpenStudy (wolf1728):

thanks :-)